# Towards Generalizable Vision-Language Robotic Manipulation: A Benchmark and LLM-guided 3D Policy

### v1

@@ V-A Experimental setup
3D-LOTUS 81.5±2.0 2.1 2.2333We use 1 V100 GPU for training for fair comparison of training time.
@@ V-D Real world experiments
Experimental setup. Our real robot setup includes a 6-DoF UR5 robotic arm equipped with three RealSense d435 cameras. We consider 7 variations across 5 tasks during training: stack cup (yellow in pink or navy in yellow), put fruit (strawberry or peach) in box, open drawer, put item in drawer and hang mug. For each task variation, we collect 20 human demonstrations via tele-operation. Then, we evaluate on the same 7 seen task variations with different objects placements and evaluate generalization capabilities on 4 new unseen task variations: put fruit (lemon and banana) in box, put food (plum then corn) in box and put fruits in plates (grapes in the yellow plate and banana in the pink plate. These tasks are illustrated in Figure 4. For each task variation we run models 10 times and report the success rate.
Table VI shows that our new method, 3D-LOTUS, outperforms PolarNet, achieving an average success rate of $8.1/10$ compared to PolarNet’s average performance of $6.7/10$. However, when applying the same 3D-LOTUS model to new task variations, we observe a complete failure to generalize to the new objects and instructions as shown in Table VII. In contrast, our improved model, 3D-LOTUS++, successfully addresses these new task variations, achieving an average success rate of $8.3/10$.
@@
Put plum in box, then corn in box 0/10 8/10
@@ VI Conclusion
ACKNOWLEDGMENT
This work was partially supported by the HPC resources from GENCI-IDRIS (Grant 20XX-AD011012122 and AD011014846).
@@ VI-C1 Task planning
Figure 6 illustrates the prompts used in LLMs for task planning. For each task variation in the training set, we craft a corresponding example as shown in Figure 7. During inference, for each new instruction, we use SentenceBert [53] to compute the sentence embedding of the instruction and compare it to the existing instructions. We then select the top 20 examples with the highest similarities to the query instruction as in-context examples to the LLM.
@@ VI-E Detail task specification
Task: Pick up the cube and put it inside the safe on the specified shelf. The shelf has three placement locations: top, middle, bottom..
@@
Success Metric: The revolute joint of the toilet seat is in the open configuratlion.
@@
Success Metric: The cubes are stacked in the correct order and on top of the green platform. .
@@
Task: Put all the .

### v2

@@ V-A Experimental setup
3D-LOTUS 83.1±0.8 2.2 2.2333We use 1 V100 GPU for training for fair comparison of training time.
@@ V-D Real world experiments
Experimental setup. Our real robot setup includes a 6-DoF UR5 robotic arm equipped with three RealSense d435 cameras. We consider 7 variations across 5 tasks during training: stack cup (yellow in pink or navy in yellow), put fruit (strawberry or peach) in box, open drawer, put item in drawer and hang mug. For each task variation, we collect 20 human demonstrations via tele-operation. Then, we evaluate on the same 7 seen task variations with different objects placements and evaluate generalization capabilities on 7 new unseen task variations: put fruit (lemon and banana) in box, put food (tuna can then corn) in box and put fruits in plates (grapes in the yellow plate and banana in the pink plate). These tasks are illustrated in Figure 4. For each task variation we run models 10 times and report the success rate.
Table VI shows that our new method, 3D-LOTUS, outperforms PolarNet, achieving an average success rate of $8.1/10$ compared to PolarNet’s average performance of $6.7/10$. However, when applying the same 3D-LOTUS model to new task variations, we observe a complete failure to generalize to the new objects and instructions as shown in Table VII. In contrast, our improved model, 3D-LOTUS++, successfully addresses these new task variations, achieving an average success rate of $7.9/10$.
@@
Stack red cup in yellow cup 0/10 8/10
Stack black cup in orange cup 0/10 7/10
Place the yellow cup inside the red cup,
then the cyan cup on top 0/10 7/10
@@
Put tuna can in box, then corn in box 0/10 8/10
@@ VI Conclusion
Acknowledgements. This work was partially supported by the HPC resources from GENCI-IDRIS (Grant 20XX-AD011012122 and AD011014846). It was funded in part by the French government under management of Agence Nationale de la Recherche as part of the “France 2030” program, reference ANR-23-IACL-0008 (PR[AI]RIE-PSAI projet), and the ANR project VideoPredict (ANR-21-FAI1-0002-01). Cordelia Schmid would like to acknowledge the support by the Körber European Science Prize.
@@ VI-C1 Task planning
Figure 6 illustrates the prompts used in LLMs for task planning. For each task variation in the training set, we craft a corresponding example as shown in Figure 7. During inference, for each new instruction, we use SentenceBert [reimers2019sentencebert] to compute the sentence embedding of the instruction and compare it to the existing instructions. We then select the top 20 examples with the highest similarities to the query instruction as in-context examples to the LLM.
@@ VI-E Detail task specification
Success Metric: The revolute joint of the fridge door has been rotated so that the door is closed and in contact with the fridge cabinet.
@@
Success Metric: The revolute joint of the laptop lid has been rotated so that the laptop is closed.
@@
Success Metric: The revolute joint of the microwave has been rotated so that the microwave door is closed.
@@
Task: Pick up the cube and put it inside the safe on the specified shelf. The shelf has three placement locations: top, middle, bottom.
@@
Success Metric: The revolute joint of the toilet seat is in the open configuration.
@@
Success Metric: The cubes are stacked in the correct order and on top of the green platform.
@@
Task: Grab each one of the groceries and put them inside the cupboard until all of them are inside the cupboard.